# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

### v2

@@ 3.4.2. Assignment of Weights: Finding the Weight Ranges
Our team collaborated with the agent creator (Team Game’s lead) to assign weights for the mental model rubric. Team Game’s lead finalized specific weights for certain concepts and a range of weights for others (Table 6). A concept with a range of weights indicates that any values within the range would be appropriate for the mental model rubric, as defined by Team Game’s lead. Our team also provided feedback on possible weights when Team Game’s lead made the decisions. This collaboration ensured our research team and the AI developers both were well-informed about the final weights, aligning with the shared goals of developing the mental model rubric.
From the weight ranges assigned by Team Game’s lead, we got four categories of weight ranges (Table 8). We labeled these categories based on how Team Game’s lead intended the weights to map the importance of the concepts. From these, we derived Table 8’s final weights as follows.
Table 7. Weight categories
Category Weight Range
Low Importance [1, 2], [-2,-1]
Low-Medium Importance [3, 5], [-5, -3]
Medium Importance [6, 10]
High Importance [11, 15]
Table 8. Criteria for weight selection
Criterion Flaws *HAB (Correct) HAB (INcorrect)
@@
Criterion 3: INcorrect-Effect-check Max Min Middle
Criterion 4: Flaw-Effect-check Min Max Max
*HAB = HowAgentBehaves
@@ 3.4.3. Assignment of Weights: Finalizing the weights
To finalize the weights, we came up with four weighting strategies, described via Table 8’s four criteria. These criteria analyzed the possible effects of the two concept groups within the mental model rubric: i) Flaws, ii) HowAgentBehaves (Correct), and HowAgentBehaves (INcorrect). To check the effects of these concepts, we have considered different combinations of weights (min, middle, max).
For Criteria 1 and 2, we assigned the maximum and minimum weights respectively within their ranges. Criterion 3 varies only in INcorrect weights from Criterion 2 to check if “INcorrect” weight makes much difference. Criterion 4: Flaw-check: varies only Flaws’ weight from Criterion 1 to check if “Flaws” weight makes much difference. As an example of the Min, Middle, and Max weights—the concept “Square X is miscalculated” has a weight range of [6, 10], within this range, we considered its min weight as 6/7, middle weight as 8, and max weight as 9/10.
The rationale behind the weights in the columns of Table 8 were:
For the Flaws concepts, the agent creator emphasized the importance of detecting the flaws as it highlights a participant’s understanding of the agent’s gameplay. However, the agent creator also noted that detecting the flaws was not very difficult, which is why we assigned lower values from their respective weight ranges.
The agent creator reasoned that HowAgentBehaves(correct) provides critical insights into the participant’s overall understanding of the agent’s behavior and recommended assigning higher weights for these concepts. So, we assigned the maximum possible values within their respective weight range in all four criteria.
Lastly, the agent creator wanted the HowAgentBehaves (INcorrect) concepts to account for participants’ incorrect understanding of the AI agents. To maximize the penalty for incorrect understanding of the AI agents, we assigned Max values to these concepts within their ranges.
Table 9 shows the four sets of weights using Table 8’s four criteria/weighting strategies.
Table 9. Exact concept weights used in the four criteria. All are within the ranges specified in Table 6, and were chosen as per the policies in Table 8. C-1 = Criterion 1, C-2 = Criterion 2 etc.
@@
Square X is miscalculated 8 7 7 7
Agent miss selection 7 7 7 7
@@
4. RQ1 Results: Which participants had better mental model scores, and why?
To answer RQ1, we needed to calculate each participant’s mental score from the concept weights and concept scores, using one of the four criteria from Section 3.1.3’s Table 9—but which one? Fortunately, it turned out that the choice among these four criteria made little difference: the four criteria generated similar participant rankings (Figure 6). For example, 9/10 of the top ten participants by Criterion 1 are also in the top ten participants of Criteria 2, 3, and 4. Likewise, 10/10 of the bottom ten are identical in all four criteria. Indeed, Criteria 2,3, and 4 give almost identical ranking results, as their almost identical overlapping lines show.
Figure 6. Mental model scores (y-axis) for each participant (x-axis) using Criterion 1 (blue) vs. Criterion 2 (red) vs. Criterion 3 (orange) vs. Criterion 4 (purple). Participants are ordered by rank (highest to lowest mental model scores as per Criterion 1). Criteria 2, 3, and 4 rank participants almost identically, and all three rank them similarly to Criterion 1.
We rejected Criterion 1 because its line graph was a little different from the other three criteria, which closely agreed with one another (Figure 6). Of the remaining three, we selected Criterion 4: Flaw-Check for two reasons: first, it avoids overemphasizing “Flaws” concepts and second, because it highlights “HowAgentBehaves” concepts. “Flaws” responses came from the games’ progressing stages, when participants’ mental models were still evolving, so seemed less likely to reflect participants’ final understanding. In contrast, “HowAgentBehaves” responses coming from the end-of-game AAR/AI questions, represented participants’ final mental models by the end of the task, so we assigned maximum weight to the “HowAgentBehaves” concepts. From here on, the term “mental model scores” will refer to the scores calculated using Criterion 4.
4.1. Which participants had better mental model scores?
To compare the two participant groups’ mental model scores, we used a one-tailed t-test. One-tailed t-tests are suitable for detecting effects occurring only in one direction (Frost, 2021), whereas two-tailed tests are for both positive and negative effects. The literature points to one direction only, unanimously reporting GenderMag to improve the user experience of people with diverse problem-solving styles  (Guizani et al., 2022; Vorvoreanu et al., 2019; Burnett et al., 2017; Hilderbrand et al., 2020; Blackmon et al., 2003; Green et al., 2000). Thus, we used Student’s one-tailed t-test to test the following hypotheses, using the mental model score calculations described in Section 3.4.3:
  • Null Hypothesis (H0): The Post-GenderMag participants’ mental model scores are less than or equal to the Original participants’ mental model scores.
  • Alternative Hypothesis (H1): The Post-GenderMag participants’ mental model scores are higher than the Original participants’ mental model scores.
@@ 4.2. Why?—Three Concepts
Why did Post-GenderMag participants develop a better mental model than the Original participants? To find out, we compared each group’s writings about the mental model concepts defined in Section 3.1.3 to identify the key concepts that differentiated the Post-GenderMag participants from the Original participants.
Figure 7 illustrates the differences between the two groups’ written comments about the mental model concepts: concepts above the dashed line were correct statements about how this particular AI worked, and concepts below the dashed line were incorrect statements. (Recall the complete definitions of each concept from Table 6.)
Figure 7. Percentage of Original(n=34) and Post-GenderMag(n=35) participants whose responses included >=>=> =1 statement about that concept. Correct concepts (above the dashed line) increase the mental model score, and incorrect concepts decrease it.
@@
Three concepts particularly differentiated the Post-GenderMag participants’ remarks from those of the Original participants: “Agent picks high score”, “Agent doesn’t consider opponent’s moves,” and “Agent is random/illogical”. To measure how much each concept differentiated the two groups, we calculated the odds of Post-GenderMag participants vs. Original participants expressing these concepts; the odds ratios are also in Figure 7. For example, as the figure shows, Post-GenderMag participants were 4.6 times more likely to correctly say that the “Agent doesn’t consider opponent’s moves”; 4 times more likely to correctly say that the “Agent picks high score”; and $\sim$ 4 times less likely to incorrectly say that the “Agent is random/illogical” than the Original participants.
@@ 4.3. The fixes behind the differentiating concepts
These significant differences in participants’ acquisition of the three differentiating concepts raise the question of what particular fixes were tied to these particular differences? We again turned to Post-GenderMag participants’ written comments to see what they had to say about this.
@@
For this purpose, we considered only responses to questions about explanations, because 9 of Team Game’s 13 fixes were changes to explanations. Within those responses, two coders independently coded Post-GenderMag participants’ references to each fix. Their inter-rater reliability (IRR) was 83% with 20% of the data. Given this acceptable level of consensus (Stemler, 2004), the coders then split up the coding of the rest of the data. Figure 8 shows the resulting counts for each fix. As the figure shows, two fixes stood out—Fix-2 and Fix-7.
@@
Figure 8. Total number of times Post-GenderMagparticipants referred to each fix. Fix-2 and Fix-7 stood out—at least 25% of the Post-GenderMag participants referred to these two fixes.
4.3.1. Fix-2: Providing exact win/loss%
In Fix-2 (Figure 3-Bottom), Team Game addressed the issue of incomplete information in the explanations, by providing a detailed breakdown of Agent Blue’s decisions. By highlighting the agent’s predicted win, loss, and draw probabilities, this fix promoted all three differentiating concepts (the agent picking from its own high scores, not using the opponent’s scores, not randomly selecting).
Specifically, whereas in the Original version (Figure 3: Top), participants could only guess at an agent’s exact scoring (win%-loss%) using the graph’s placement of the tiny rectangle of interest, Fix-2 added the option to see the exact win/loss %s with each rectangle of the explanations (Figure  3: Bottom). Upon hovering over a rectangle, the participants could see the exact win%, loss%, and drawif Agent Blue picked the square represented by this rectangle.
@@ 4.3.1. Fix-2: Providing exact win/loss%
Viewing these exact win/loss/draw %s may explain why Post-GenderMag participants were four times more likely to notice “Agent picks high score” (Figure 7). As one Post-GenderMag participant wrote:
@@
Figure 9. Top: (a) Before Fix-2 (Original version): Hovering over a rectangle from the latest data series highlights the rectangles representing the same gameboard square in every data series [^3] Every time agent blue makes a move, the BTW explanation generates a data series containing 36 small rectangles, 1 rectangle for each gameboard square. Where the rectangle lies vertically tells how good the agent thinks moving that square would be. Multiple data series show this info for different decision points.from the previous decision points, but does not show the exact probabilities and does not show which rectangles came from which data series.
Bottom: (b) After Fix-2 (Post-GenderMag version): Hovering over the rectangle now explicitly shows the win%, loss%, and drawof the gameboard square represented by that rectangle.
@@
Fix-2’s emphasis on the agent’s own calculations also may have helped most Post-GenderMag participants avoid an error that the Original participants made: thinking that the “Agent is random/illogical”, which is not true. Compared to Post-GenderMag participants, Original participants were 4 times more likely to incorrectly assume that Agent Blue was random/illogical. In contrast, Post-GenderMag participants’ responses reflected an understanding of the logic behind Agent Blue’s decisions:
P02-PostGM: BTW allowed me to assess all possibilities the AI was considering & how it percieved it’s win % ¡Available with Fix-2¿.
@@
4.3.2. Fix-7: Provided move ranking
Besides Fix-2’s exact win/loss%, Team Game decided to enable participants to compare the top-ranked squares. So with Fix-7, they added an explanation “Top 5 Moves”, which they implemented via Fix-7 (Figure 10). Unlike Fix-2, Fix-7 provided the exact score, which emphasized that Agent Blue usually followed its scores and positions among the top scores. Seeing the exact scores and comparing the top-ranked moves could be the reasons why more Post-GenderMag participants said that “Agent picks high score”:
@@ 4.3.2. Fix-7: Provided move ranking
Participant P39-PostGM also explained how the top 5 moves helped the participant see the agent’s behavior as being purposeful, not random/illogical:
@@
4.3.3. Causes and the Effects of the fixes
Causes: Both Fix-2 and Fix-7 were about adding more information, driven by insights from Team Game’s GenderMag analyses. As Figure 11 shows Team Game kept pointing to the need for more information because the facet Information Processing Style dominated the GenderMag sessions with Motivations and Attitude Toward Risk following closely behind.
Figure 11. The number of times Team Game mentioned the GenderMag facets during the GenderMag sessions that led to the development of Fix-2 and Fix-7.
@@ 4.3.3. Causes and the Effects of the fixes
5. RQ2 Results: Prediction Performance and why
5.1. The Prediction Results
To triangulate a different evidence source with the mental model, next, we compared participants’ prediction performance. We calculate participants’ prediction performance using Dodge et al.’s method of calculating “Loss in Value” each per prediction (Koujalgi et al., 2024). Loss in value is a measure of how erroneous a prediction is, we term it PredError in this paper.
PredError’s calculation is:

### v3

@@ 3.6.2. Assignment of Weights: Finding the Weight Ranges
Team Game’s lead finalized specific weights for certain concepts and a range of weights for others (Table 6). A concept with a range of weights indicates that any values within the range would be appropriate for the mental model concepts rubric, as defined by Team Game’s lead. Our team also provided feedback on possible weights when Team Game’s lead made the decisions. From these, we derived Table 7’s final weights as follows.
Table 7. Criteria for weight selection derived from Table 6’s weight ranges. HAB = How Agent Behaves.
Criterion Flaws HAB (Correct) HAB (INcorrect)
@@
Criterion 3: INcorrects’ effect check Min Max Middle
Criterion 4: Flaws’ effect check Min Max Max
@@ 3.6.3. Assignment of Weights: Finalizing the weights
To finalize the weights, we came up with four weighting strategies, described via Table 7’s four criteria. These criteria analyzed the possible effects of the two concept code groups within the mental model concepts rubric: (i) Flaws, (ii) How Agent Behaves (Correct), and How Agent Behaves (INcorrect). To check the effects of these groups, we considered different combinations of weights (min, middle, max).
For Criteria 1 and 2, we assigned all the concepts with maximum/middle and minimum weights within their ranges respectively. Criterion 3 varies from Criterion 2 only in INcorrect concepts’ weights to check if “INcorrect” concepts make much difference within the mental model concepts rubric. Lastly, Criterion 4 varies from Criterion 1 only in Flaws concepts’ weights to check if “Flaws” concepts make much difference. As an example of the Min, Middle, and Max weights is the concept “Square ‘X’ is miscalculated,” which has a weight range of [6, 10]. Within this range, we considered its min weight as 7, middle weight as 8, and max weight as 10.
The reasonings behind the weights in the columns of Table 7 were as follows.
For the Flaws concepts, the agent creator emphasized the importance of detecting the flaws as it shows a participant’s understanding of the agent’s gameplay. However, the agent creator also noted that detecting the flaws was not very difficult. Thus, we assigned low weights to the flaws’ concepts with “min” weights in three criteria, “middle” weights in one, and “max” weights in none.
Next, the agent creator concluded that How Agent Behaves (correct) concepts provide critical insights into the participant’s overall understanding of the agent’s behavior and recommended assigning higher weights for these concepts. So, we assigned the maximum possible weights within these concepts’ ranges in all four criteria.
Lastly, the agent creator wanted the How Agent Behaves (INcorrect) concepts to account for participants’ incorrect understanding of the AI agents. However, they didn’t specify any priorities for these concepts. So, we tested all three—min, middle, and max weights for these concepts within the four criteria.
Table 8 shows all the concepts’ exact weights that we derived from Table 7’s four criteria.
Table 8. Exact concept weights used in the four criteria. All are within the ranges specified in Table 6, and were chosen as per the strategies in Table 7. C-1 = Criterion 1, C-2 = Criterion 2 etc.
@@
Square “X” is miscalculated 8 7 7 7
Agent misselection 7 7 7 7
@@
4. RQ1 Results: Which Participants Had Better Mental Model Concepts Scores, and Why?
Curb-cut effects are a pair of effects—a positive “better for everyone” effect and a “better for an underserved group” effect. We begin by considering the “better for everyone” question.
To answer RQ1’s question about participants’ mental model concepts, we needed to calculate each participant’s mental model concepts scores using the concept weights and concept counts [^1] The concept count is the total number of times we identified a particular concept in a participant’s responses., using one of the four criteria from Table 7—but which one? Fortunately, it turned out that the choice among these four criteria made little difference: the four criteria generated similar participant rankings (Figure 6). For example, 9/10 of the top ten participants by Criterion 1 were also the top ten participants of Criteria 2, 3, and 4. Likewise, 10/10 of the bottom ten were identical in all four criteria. Indeed, Criteria 2, 3, and 4 gave almost identical ranking results, as their almost identical overlapping lines show (Figure 6).
Figure 6. Mental model concepts scores (y-axis) for each participant (x-axis) using Criterion 1 (blue) vs. Criterion 2 (red) vs. Criterion 3 (orange) vs. Criterion 4 (purple). Participants are ordered by rank (highest to lowest mental model concepts scores as per Criterion 1). Criteria 2, 3, and 4 rank participants almost identically, and all three rank them similarly to Criterion 1.
In choosing one among such similarly behaving criteria, we rejected Criterion 1 because its line graph was a little different from the other three criteria. Of the remaining three, we selected Criterion 4: Flaws’ effect check for two reasons. First, it avoids overemphasizing the “Flaws” concepts. This was important because “Flaws” responses came throughout the games’ progressing stages, when participants’ mental models were still evolving, so sometimes it did not reflect participants’ final understanding. Second, Criterion 4 emphasizes “How Agent Behaves” concepts, which came from the end-of-game AAR/AI questions and thus represented participants’ final mental models. From here on, the term “mental model concepts scores” will refer to the scores calculated using Criterion 4.
4.1. Which Participants Had Better Mental Model Concepts Scores?
@@ 4.2. Why?—Three Concepts
To find out why Post-GenderMag participants had higher mental model concepts scores than the Original participants, we compared each group’s responses about the mental model concepts defined in Section 3.6.1 to identify which ones differentiated the Post-GenderMag participants from the Original participants.
Figure 7 shows the results: concepts above the dashed line were correct statements about how this particular AI worked, and concepts below the dashed line were incorrect statements. (Recall the complete definitions of each concept from Table 6.)
Figure 7. Percentage of Original(n=34) and Post-GenderMag(n=35) participants whose responses included >=>=> =1 statement about that concept. Correct concepts (above the dashed line) increase the mental model concepts score, and incorrect concepts decrease it.
@@
Three concepts particularly differentiated the Post-GenderMag participants’ remarks from those of the Original participants: “Agent picks high score”, “Agent doesn’t consider opponent’s moves,” and “Agent is random/illogical”. To measure how much each concept differentiated the two groups, we calculated the odds of Post-GenderMag participants vs. Original participants expressing these concepts; the odds ratios are shown in Figure 7. For example, as the figure shows, Post-GenderMag participants were 4.56 times more likely to correctly say that the “Agent doesn’t consider opponent’s moves”; 4.04 times more likely to correctly say that the “Agent picks high score”; and 3.7 ($1/0.27$) times less likely to incorrectly say that the “Agent is random/illogical” than the Original participants.
@@ 4.3. The Fixes Behind the Differentiating Concepts
These significant differences in participants’ acquisition of the three differentiating concepts raise the question of what particular fixes were tied to these particular differences. We again turned to Post-GenderMag participants’ responses to see what they had to say about this.
@@
For this purpose, we considered only responses to questions about explanations (explanation feedback in Table 3), because 9 of Team Game’s 13 fixes were changes to explanations. Within those responses, two coders independently coded Post-GenderMag participants’ references to each fix. Their inter-rater reliability (IRR) was 83% with 20% of the data. Given this acceptable level of consensus (Stemler, 2004), the coders then split up the coding of the rest of the data. Figure 8 shows the resulting counts for each fix. As the figure shows, two fixes stood out—Fix-2 and Fix-7.
@@
Figure 8. Total number of times Post-GenderMag participants referred to each fix. Another quantitative way (not shown) Fix-2 and Fix-7 stood out was that 26/35 (74%) and 12/35 (34%) individual participants referred to Fix-2 and Fix-7, respectively.
Recall that these fixes were triggered by Team Game’s earlier analysis of the “inclusivity bugs” in their Original prototype, and were then designed by Team Game to resolve the bugs in whatever way they saw fit (Anderson, 2024). During that inclusivity analysis, Team Game had noticed several instances of leaving out users whose Information Processing Style was the “comprehensive” style; i.e., those who like to gather a sizable batch of relevant information before proceeding with a problem-solving action or decision. In fact, as Figure 9 shows, Team Game talked about this problem-solving style more than any other during their discussions leading to these fixes. Ultimately, they devised Fix-2 and Fix-7 to add more information in order to be more inclusive to users with comprehensive Information Processing Styles, who had not been explicitly considered in the Original prototype.
Figure 9. Number of times Team Game mentioned the GenderMag problem-solving styles during the GenderMag sessions that led to the development of Fix-2 and Fix-7.
[…]
4.3.1. Fix-2: Exact win/lossavailable
To increase support across the range of diverse users’ Information Processing Styles while also addressing diverse Learning Styles, Motivations, and Attitudes Towards Risk, in Fix-2 (Figure 4: Bottom) Team Game created a detailed breakdown of Agent Blue’s decisions, available via a mouse hover. Specifically, whereas in the Original version (Figure 4: Top), participants could only guess at an agent’s exact scoring (win%-loss%) using the graph’s placement of the tiny rectangle of interest, Fix-2 enabled participants to hover over an explanation’s rectangle to see the exact win%, loss%, and drawif Agent Blue picked the move represented by this rectangle (Figure 4: Bottom). Note how this feature supports a wide range of Information Processing Styles: comprehensive information processors can hover to see it, and selective information processors can simply not hover. Perhaps this is why it was so well-received: 74% of the Post-GenderMag participants referred to it at some point during the study.
Figure 10. Top: (a) Before Fix-2 (Original version). Hovering over a rectangle from the latest data series highlights the rectangles representing the same gameboard square in every data series [^4] Every time Agent Blue makes a move, the BestToWorst explanation generates a data series containing 36 small rectangles, 1 rectangle for each gameboard square. Where the rectangle lies vertically tells how good the agent thinks moving that square would be. Multiple data series show this info for different decision points. from the previous decision points, but does not show the exact probabilities and does not show which rectangles came from which data series.
Bottom: (b) After Fix-2 (Post-GenderMag version). Hovering over the rectangle now explicitly shows the win%, loss%, and drawof the gameboard square represented by that rectangle.
@@ 4.3.1. Fix-2: Exact win/lossavailable
The ability to view these exact win/loss/draw %s may explain why Post-GenderMag participants were $4.04$ times more likely to notice “Agent picks high score”. As one Post-GenderMag participant wrote:
@@
Fix-2’s emphasis on the agent’s own calculations also may have helped most Post-GenderMag participants avoid an error that the Original participants made: thinking that the “Agent is random/illogical”, which is not true. Compared to Post-GenderMag participants, Original participants were 3.7 ($1/0.27$) times more likely to incorrectly assume that Agent Blue was random/illogical. In contrast, Post-GenderMag participants’ responses reflected an understanding of the logic behind Agent Blue’s decisions:
P02-PostGM: BTW ¡BestToWorst¿ allowed me to assess all possibilities the AI was considering & how it percieved it’s win % ¡Available with Fix-2¿.
@@
4.3.2. Fix-7: Add move rankings for comparison
Finding similarities and differences is a central component of many cognitive processes  (Markman and Gentner, 1996), and Team Game decided that supporting this ability would be particularly helpful to comprehensive information processors. Thus, with Fix-7, Team Game added a new explanation “Top 5 Moves” (Figure 11), which enabled participants to compare the scores and win/loss/draw %s of the top contenders for Agent Blue’s prior move.
Seeing the exact scores and comparing the top-ranked moves could be the reasons why more Post-GenderMag than Original participants said that “Agent picks high score”:
@@ 4.3.2. Fix-7: Add move rankings for comparison
Participant P39-PostGM also explained how the top 5 moves helped them see the agent’s behavior as being purposeful, not random/illogical:
@@
Note a nuance of Fix-7: it follows a different inclusivity strategy than Fix-2 did. With Fix-2, participants who leaned toward a comprehensive Information Processing Style had to mouse hover to retrieve the additional information Fix-2 provided. Thus, participants who did not want more information were not “force-fed” it. In contrast, Fix-7’s additional information was always on-screen. This still does not force everyone to read it—arguably, a user can always ignore it—but its constant presence communicates an encouragement for everyone to read it. In Section 5.2, we will return to ways Fix-7 may have affected participants’ mental models.
4.4. The Post-GenderMag Participants’ Mental Models and Curb Cuts
In the Introduction, we alluded to the “curb cut effect,” which occurs when using inclusive design to improve an environment for underserved users helps everyone—including those who were not previously underserved. The results just presented suggests that this is exactly what happened in this study. Team Game’s inclusivity fixes to the Post-GenderMag prototype’s explanations, which they designed to make their explanations more inclusive for “Abi-like” users (Section 2.1), actually led to significantly higher overall mental model concepts scores for the entire Post-GenderMag group than in the Original group.
@@ 4.4. The Post-GenderMag Participants’ Mental Models and Curb Cuts
However, the mental model concepts scoring used in this section is not the only way to measure mental models. We next consider what a different measure of participants’ mental models reveals.
5. RQ2 Results: Prediction Accuracy and Why
5.1. The Prediction Accuracy Results
The mental model concepts score used in the previous section measured participants’ mental models in terms of concepts of the AI’s behavior that participants understood. In this section, we consider an alternative measure of mental models, namely participants’ ability to predict the AI’s next behavior. Whereas the mental model concepts score measured participants’ conceptual understanding of the AI’s behaviors, measuring a participant’s prediction accuracy measures a participant’s ability to actually mimic what the AI does, namely choose the same next behavior.
We calculate participants’ prediction accuracy using the absolute value of Dodge et al.’s method of calculating “Loss in Value” per prediction (Koujalgi et al., 2024). Loss in value is a measure of how erroneous a prediction is; in this paper, we term it PredError. Its calculation is: